Using dyplr for R:

Create data frame where each student is represented once.

Example of what I have:

student test score
1 1 4
1 2 5
1 3 6

What it should look like:

Student score 1 score 2 score 3 score 4
1 4 5 6 8
2 1 5 6 8
3 4 3 6 9
4 2 3 4 5
Answer #1

Solution:

Student = c(1, 2, 3, 4)
score1 = c(4, 1, 4, 2)
score2 = c(5, 5, 3, 3)
score3 = c(6, 6, 6, 4)
score4 = c(8, 8, 9, 5)

Students = data.frame(Student, score1, score2, score3, score4)
print(Students)

Output: